Marlin24k:

Solve: 7x+9+2x+90=180 solve for x

dude:

Isolate x first

Joe348:

Combine the like terms on the left side of equation \[7x+2x=9x\] \[9+90=99\] So the left side of the equation is \[9x+99\] \[9x+99=180\] Subtract 99 from both sides \[9x=81\] Divide both sides by 9 \[x= 81 \div 9= \]